OpenStudy (anonymous):

(x+3)^1/2-1=x I know the answer is 1 but do not know how to get it.

hero (hero):

\(\sqrt{(x + 3)} - 1 = x\) Add 1 to both sides: \(\sqrt{(x + 3)}= x + 1\) Square both sides: \(x + 3 = (x + 1)^2\) Expand the right side, then put in the form \(ax^2 + bx + c = 0\)